OpenStudy (anonymous):

A jar contains 8 green, 5 red, 4 blue, and 3 black marbles.once a marble is selected, it is not replaced. find each probability. 1) p(two green marbles) 2) p(two black marbles) 3) p(a red marble and then a blue marble) 4) p(a black marble then a green marble) 5) p(two marbles that are not red) 6) p(two marbles that are neither blue nor green) 7) a family has two children, both of whom are boys. what is the probability that the next child will be born a boy?

OpenStudy (anonymous):

how about one at a time?

OpenStudy (anonymous):

\[\frac{8}{20}\times \frac{7}{19}\] for the first one